OpenStudy (helpblahblahblah):

MEDAL: Four friends went scuba diving. According to the following statements, who spotted sea life at the fastest rate? Lance saw 2 reef sharks in 6 minutes. Jasmine saw 3 sea cucumbers in 8 minutes. Robert saw 4 jellyfish in 15 minutes. Shilpa saw 3 pencil urchins in 10 minutes.

OpenStudy (anonymous):

2 divided by 6 = ? 3 / 8 = ? 4 / 15 = ? 3 / 10 = ? then compare

OpenStudy (helpblahblahblah):

thanks! :D

OpenStudy (anonymous):

That's exactly what I was about to say.